OpenStudy (anonymous):

What are the odds of rolling a 6 on a regular six sided die?

OpenStudy (anonymous):

@Michele_Laino

OpenStudy (anonymous):

i think its 1:6 @Michele_Laino

OpenStudy (michele_laino):

I think that they are 1, 3, and 5

OpenStudy (michele_laino):

namely the odd numbers are 1, 3, and 5

OpenStudy (michele_laino):

the requested probability is: \[\frac{{{\text{favorable outcomes}}}}{{{\text{possible outcomes}}}} = \frac{1}{6}\]

OpenStudy (anonymous):

IS ODDS DIFFERENT THE PROBABILITY @Michele_Laino

OpenStudy (anonymous):

1/6

OpenStudy (anonymous):

but isnt probability and odds different

OpenStudy (michele_laino):

for example, the odds is the probability when all the possible outcomes are different each from other

OpenStudy (anonymous):

so 1:6 would be correct?

OpenStudy (anonymous):

for odds?

OpenStudy (michele_laino):

I'm pondering

OpenStudy (anonymous):

@Odds is defined as success : failures. There are two numbers greater than or equal to 5. Those numbers are 5 and 6. The failures would be the numbers that are not greater than or equal to 5. Those numbers are 1, 2, 3, and 4. This means that odds is 2 : 4 or 1 : 2.

OpenStudy (anonymous):

im confused.

OpenStudy (anonymous):

probabily would be 1:6 but for odds it would be 1:5 correct?

OpenStudy (michele_laino):

here is the right definition of odds: "odds is the ratio between the probability that an event occurs and the probability that the same event not occurs" so for question #1 probability to get "m" = 1/7 probability to get not an "m" = 1-(1/7)=6/7 \[\Large odds = \frac{{1/7}}{{6/7}} = \frac{1}{6}\]

OpenStudy (michele_laino):

another example: for a six sided dice, we have: probability to get a "6" = 1/6 probability to get not a "6" = 1-(1/6)= 5/6 so: \[\Large odds = \frac{{1/6}}{{5/6}} = \frac{1}{5}\]

OpenStudy (michele_laino):

finally: question #3 probability to get "g" = 2/7 probability to get not a "g" = 1-(2/7) = 5/7 so: \[\Large odds = \frac{{2/7}}{{5/7}} = \frac{2}{5}\]

OpenStudy (anonymous):

Such a complicated process to find out something so simple! Would've never guessed. Thanks for broadening my horizons :) @Michele_Laino

OpenStudy (michele_laino):

:)

OpenStudy (jack1):

1 chance in 6 possibilities ;)

OpenStudy (anonymous):

1/6